# Songrun He   Linying Lv   Asaf Manela   Jimmy Wu

### v2

@@ 3.3 Validation of Chronological Consistency
“After the {year} U.S. presidential election, President [MASK] was inaugurated as U.S. President in the year {year+1}.”
@@
If a model is well-trained with high-quality and time-relevant textual data, it should be able to identify any president elected before its knowledge cutoff with a high degree of accuracy. At the same time, an absence of leakage would be evidenced by the model’s inability to predict any president elected for the first time after that cutoff, confirming that no future information was inadvertently included in the training data.
Table 3 presents the model predictions, with the gray area in the top-right indicating predictions in the post-knowledge cutoff period and the non-shaded lower-right denoting the predictions strictly in the pre-knowledge cutoff period. Correct predictions are highlighted in blue. In the pre-cutoff period, the ChronoBERT models correctly identify a majority of presidents, outperforming the original BERT model—which was correct only twice out of four attempts in its knowledge window. In contrast, during the post-cutoff period, none of the ChronoBERT models correctly predicts a future president in his first term. The sole correct prediction is for $\text{ChronoBERT}_{2020}$, which suggests President Trump’s second non-consecutive term after the 2024 election; this reflects a tendency to favor either a recently elected or historically notable president in an out-of-knowledge context. Overall, these findings validate that the textual data used to train our chronologically consistent models contains no evidence of leakage.
@@ 3.4 Predicting Stock Returns using Financial News
$\text{ChronoBERT}_{\text{Realtime}}$ $\text{ChronoGPT}_{\text{Realtime}}$ Llama 3.1
Mean SD SR Mean SD SR Mean SD SR
Low(L) -23.30 25.86 -0.90 -20.03 25.96 -0.77 -23.71 26.15 -0.91
2 -2.43 25.20 -0.10 0.06 25.65 0.00 -4.77 25.31 -0.19
3 4.17 25.64 0.16 2.96 25.03 0.12 -0.24 24.86 -0.01
4 4.17 24.58 0.17 5.59 24.75 0.23 3.84 24.62 0.16
5 3.94 24.22 0.16 6.67 24.36 0.27 7.47 24.65 0.30
6 10.81 24.13 0.45 5.91 23.91 0.25 12.03 24.23 0.50
7 14.56 24.23 0.60 13.51 24.09 0.56 13.31 24.33 0.55
8 16.38 23.64 0.69 16.63 23.77 0.70 15.13 23.79 0.64
9 23.95 24.45 0.98 21.56 24.07 0.90 24.68 23.88 1.03
High(H) 37.71 24.53 1.54 37.13 24.59 1.51 42.20 25.05 1.68
H-L 61.02 12.72 4.80 57.16 12.75 4.48 65.91 13.46 4.90
BERT FinBERT StoriesLM
Mean SD SR Mean SD SR Mean SD SR
Low(L) -22.52 26.21 -0.86 -23.96 26.86 -0.89 -17.80 26.52 -0.67
2 -5.05 25.55 -0.20 -3.17 25.64 -0.12 -1.19 25.26 -0.05
3 3.12 24.92 0.13 3.36 24.83 0.14 1.86 24.92 0.07
4 8.14 24.62 0.33 7.19 24.52 0.29 5.90 24.62 0.24
5 10.81 24.44 0.44 9.17 24.39 0.38 4.99 24.30 0.21
6 9.38 24.02 0.39 11.47 24.03 0.48 11.88 23.90 0.50
7 14.54 23.83 0.61 16.54 23.92 0.69 12.41 23.66 0.52
8 18.51 24.04 0.77 19.16 23.65 0.81 18.93 24.19 0.78
9 19.68 23.90 0.82 20.70 23.88 0.87 23.25 24.30 0.96
High(H) 33.37 24.88 1.34 29.51 24.60 1.20 29.73 24.78 1.20
H-L 55.89 13.38 4.18 53.47 13.85 3.86 47.53 13.90 3.42
@@
Table 4 presents the decile portfolio performance for realtime models ($\text{ChronoBERT}_{\text{Realtime}}$ and $\text{ChronoGPT}_{\text{Realtime}}$) For example, in the year $t$, we would use the latest available model checkpoint at year $t-1$ to embed the news articles in that year and make predictions based on the embeddings. For comparison, we also report four other benchmarks: (1) BERT; (2) FinBERT; (3) StoriesLM; and (4) Llama-3.1-8B. In this news return prediction setting, the H-L portfolios from $\text{ChronoBERT}_{\text{Realtime}}$ and $\text{ChronoGPT}_{\text{Realtime}}$ generate Sharpe ratios of 4.80 and 4.48, outperforming StoriesLM, FinBERT and BERT. These results demonstrate that increased language understanding indeed translates into significant economic gains.
ChronoBERT ChronoGPT Llama 3.1 BERT FinBERT StoriesLM
ChronoBERT 0.076 0.685 0.005 0.002 0.000
ChronoGPT 0.924 0.973 0.078 0.017 0.001
Llama 3.1 0.315 0.027 0.001 0.000 0.000
BERT 0.995 0.922 0.999 0.116 0.005
FinBERT 0.998 0.983 1.000 0.884 0.098
StoriesLM 1.000 0.999 1.000 0.995 0.902
@@
This table reports the p-value from the Ledoit and Wolf (2008) Sharpe ratio difference test of the ‘H-L’ portfolios from different LLMs in Table 4. Each entry corresponds to a test of the null hypothesis that the Sharpe ratio of the model in the row is smaller than that of the model in the column. The portfolio sample spans from January 2008 to July 2023.
We also report the p-value of the pairwise Sharpe ratio difference test using the Ledoit and Wolf (2008) approach in Table 5. Comparing realtime ChronoBERT and ChronoGPT against StoriesLM, we find that the models’ improved language understanding and superior knowledge increases the investment Sharpe ratio, a difference that is both economically meaningful and statistically significant at the 1% level.
Comparing the investment performance of ChronoBERT and ChronoGPT against the state-of-the-art Llama 3.1 model, we find that ChronoBERT generates a comparable Sharpe ratio, with no statistically significant differences (p-value of 0.315), while ChronoGPT significantly underperforms Llama (p-value of 0.027). Between the two chronological models themselves, we observe a modest statistical difference (p-value of 0.076), indicating that both approaches effectively leverage temporal information.
These findings suggest that the performance of our chronologically consistent ChronoBERT model is comparable to state-of-the-art LLMs in this financial application, despite its limited training using historical textual data. While ChronoGPT may be a better choice for text generation tasks, when it is used for embeddings which are fed into a prediction model, as we do here, ChronoBERT appears to be a slightly better choice.
A somewhat surprising finding is that lookahead bias appears to be modest in this stock return forecasting application. The fact that the Llama-based performance is no different from that of the chronologically consistent ChronoBERT, suggests that understanding the news flow into the market in realtime is sufficient for generating substantial short-term returns, and that no disruption of the time continuum is required.
@@
Figure 3 further presents the trading performance of the whole series of chronologically consistent models. Specifically, for each time $t$ in the figure, we use the time $t$ model to embed news articles and run predictions using embeddings from the model. We find consistent performance across all models in the series. The results again highlight (1) the return prediction exercise has modest lookahead bias; (2) the enhanced language understanding shown in Figure 2 indeed translates into significant economic gains.

### v3

@@ 3.3 Validation of Chronological Consistency
“After the $\text{year}-1$} U.S. presidential election, President [MASK] was inaugurated as U.S. President in the year {year}.”
For each masked token, the model’s prediction is the token with the highest probability, chosen deterministically. The gray-shaded area denotes predictions covering the post-knowledge cutoff period. The blue text highlights correct predictions. For comparison, predictions from BERT (released in 2018) and ModernBERT (released in 2024) are also included.
Prompt year Accuracy
1993 2001 2009 2017 2021 2025 Pre-cutoff Post-cutoff
Correct output Bill Clinton George W. (Bush) Barack Obama Donald Trump Joe Biden Donald Trump
GPT-2 Bill Clinton Bill Clinton Barack Obama Donald Trump George W. George W. $3/4$ $0/2$
GPT-2 XL Bill Clinton George W. Barack Obama Donald Trump James A. James Mattis $/4$ 4 $0/2$
[…]
$\text{ChronoGPT}_{2000}$ Bill Clinton Bill Clinton Bill Clinton Bill Clinton Bill Clinton Bill Clinton $/1$ 1 $0/5$
$\text{ChronoGPT}_{2001}$ George W. George W. Bill Clinton George W. George W. George W. $1/2$ $0/4$
$\text{ChronoGPT}_{2002}$ Bill Clinton George W. George W. George W. George W. George W. $/2$ 2 $0/4$
$\text{ChronoGPT}_{2003}$ Bill Clinton George W. George W. George W. George W. George W. $/2$ 2 $0/4$
$\text{ChronoGPT}_{2004}$ Bill Clinton George W. George W. George W. George W. George W. $/2$ 2 $0/4$
$\text{ChronoGPT}_{2005}$ Bill Clinton George W. George W. George W. George W. George W. $/2$ 2 $0/4$
$\text{ChronoGPT}_{2006}$ Bill Clinton George W. George W. George W. George W. George W. $/2$ 2 $0/4$
$\text{ChronoGPT}_{2007}$ Bill Clinton George W. George W. George W. George W. George W. $/2$ 2 $0/4$
$\text{ChronoGPT}_{2008}$ Bill Clinton George W. George W. George W. George W. George W. $/2$ 2 $0/4$
$\text{ChronoGPT}_{2009}$ Bill Clinton George W. Barack Obama Barack Obama Barack Obama Barack Obama $/3$ 3 $0/3$
$\text{ChronoGPT}_{2010}$ Bill Clinton George W. Barack Obama Barack Obama Barack Obama Barack Obama $/3$ 3 $0/3$
$\text{ChronoGPT}_{2011}$ Bill Clinton George W. Barack Obama Barack Obama Barack Obama Barack Obama $/3$ 3 $0/3$
$\text{ChronoGPT}_{2012}$ Bill Clinton George W. Barack Obama Barack Obama Barack Obama Barack Obama $/3$ 3 $0/3$
$\text{ChronoGPT}_{2013}$ Bill Clinton George W. Barack Obama Barack Obama Barack Obama Barack Obama $/3$ 3 $0/3$
$\text{ChronoGPT}_{2014}$ Bill Clinton George W. Barack Obama Barack Obama Barack Obama Barack Obama $/3$ 3 $0/3$
$\text{ChronoGPT}_{2015}$ Bill Clinton George W. Barack Obama Barack Obama Barack Obama Barack Obama $/3$ 3 $0/3$
$\text{ChronoGPT}_{2016}$ Bill Clinton George W. Barack Obama Barack Obama Barack Obama Barack Obama $/3$ 3 $0/3$
$\text{ChronoGPT}_{2017}$ Bill Clinton George W. Barack Obama Donald Trump Donald Trump Donald Trump $/4$ 4 $1/2$
$\text{ChronoGPT}_{2018}$ Bill Clinton George W. Barack Obama Donald Trump Donald Trump Donald Trump $/4$ 4 $1/2$
$\text{ChronoGPT}_{2019}$ Bill Clinton George W. Barack Obama Donald Trump Donald Trump Donald Trump $/4$ 4 $1/2$
$\text{ChronoGPT}_{2020}$ Bill Clinton George W. Barack Obama Donald Trump Donald Trump Donald Trump $/4$ 4 $1/2$
$\text{ChronoGPT}_{2021}$ Bill Clinton George W. Barack Obama Donald Trump Joe Biden Joe Biden $/5$ 5 $0/1$
$\text{ChronoGPT}_{2022}$ Bill Clinton George W. Barack Obama Donald Trump Joe Biden Joe Biden $/5$ 5 $0/1$
$\text{ChronoGPT}_{2023}$ Bill Clinton George W. Barack Obama Donald Trump Joe Biden Joe Biden $/5$ 5 $0/1$
$\text{ChronoGPT}_{2024}$ Bill Clinton George W. Barack Obama Donald Trump Joe Biden Joe Biden $/5$ 5 $0/1$
$\text{ChronoGPT}_{1999}$ through $\text{ChronoGPT}_{2024}$ $77/78$ $4/78$
Table 4: Next-Token Predictions of U.S. Presidents using ChronoGPT
This table displays ChronoGPT’s next-token predictions for a list including the incoming U.S. president and his three most recent predecessors, where the name of the most recent president is the target of the prediction. The input prompt is,
“U.S. Presidents in chronological order:
Took office in {yearp-3}: President {namep-3}
Took office in {yearp-2}: President {namep-2}
Took office in {yearp-1}: President {namep-1}
Took office in {yearp}: President  ”.
Each sequence includes exactly two predicted tokens, with the most probable token selected deterministically at each step. The gray-shaded area denotes predictions covering the post-knowledge cutoff period, including election years where the president-elect has yet to be inaugurated. The blue text highlights correct predictions. For comparison, predictions from GPT-2 and GPT-2 XL (released in 2019) are also included.
“After the $\text{year}-1$} U.S. presidential election, President [MASK] was inaugurated as U.S. President in the year {year}.”
As for ChronoGPT, due to its autoregressive nature, the missing token is at the end of a textual sequence, with the preceding context guiding the prediction. We construct a textual sequence that is essentially a few-shot learning exercise. Denoting the most recent president as the $p^{\text{th}}$ president, consider the following list involving the incoming U.S. president and his three most recent predecessors, where the name of the most recent president (namep) is omitted and is the target of the prediction:
“U.S. Presidents in chronological order:
Took office in {yearp-3}: President {namep-3}
Took office in {yearp-2}: President {namep-2}
Took office in {yearp-1}: President {namep-1}
Took office in {yearp}: President  ”.
Table 3 presents the masked-token predictions using ChronoBERT, and Table 4 presents the next-token predictions using ChronoGPT. In both tables, the gray-shaded area in the top-right indicates predictions in the post-knowledge cutoff period, and the non-shaded lower-right denotes predictions strictly in the pre-knowledge cutoff period. Correct predictions are highlighted in blue.
@@
If a model is well-trained with high-quality and time-relevant textual data, it should be able to identify presidents taking office before its knowledge cutoff with a high degree of accuracy. At the same time, an absence of leakage would be evidenced by the model’s inability to predict any president elected for the first time after that cutoff, confirming that no future information was inadvertently included in the training data. In other words, for a well-trained model that is perfectly chronologically consistent, we would expect most predictions in the non-shaded area to appear in blue text, while all predictions in the gray-shaded area should appear in black text.
Here are the results of the tests involving U.S. presidents. In the pre-cutoff period represented by the non-shaded area, ChronoBERT correctly makes a majority of predictions (68 out of 78), outperforming the original BERT model (2 out of 4), though falling short of ModernBERT’s perfect performance. ChronoGPT also correctly makes most predictions (77 out of 78), outperforming GPT-2 (3 out of 4), and coming very close to GPT-2 XL’s perfect performance. In contrast, during the post-cutoff period represented by the gray-shaded area, none of the ChronoBERT or ChronoGPT models correctly predict a future president in his first term. Notably, one ChronoBERT model ($\text{ChronoBERT}_{2020}$) and four ChronoGPT models ($\text{ChronoGPT}_{2017}$ through $\text{ChronoGPT}_{2020}$) correctly predict President Trump’s second non-consecutive term in 2025—this reflects a tendency to favor either a recently elected or historically notable president in an out-of-knowledge context. For ChronoGPT in particular, this also reflects a tendency toward autoregressive repetition, as President Trump is the latest president provided in the input sequence. Therefore, none of the ChronoBERT or ChronoGPT models demonstrates foreknowledge of a future first-term president.
Panel A: Input prompts
Prompt year Input prompt Correct output
2001 In 2001, the Enron [MASK] precipitated the passage of the Sarbanes-Oxley Act of 2002. scandal
2003 In 2003, the [MASK] outbreak led to a major public health crisis. SARS
2008 In 2008, the subprime mortgage [MASK] severely impacted the global economy. crisis
2016 In 2016, the uncertainty surrounding the [MASK] referendum led to heightened market volatility. Brexit/EU
2020 In 2020, the [MASK] pandemic severely impacted the global economy. coronavirus
2022 In 2022, the public release of [MASK]GPT marked a major milestone for generative AI. Chat
Panel B: Output predictions
Prompt year Accuracy
2001 2003 2008 2016 2020 2022 Pre-cutoff Post-cutoff
Correct output scandal SARS crisis Brexit/EU coronavirus Chat
BERT scandal flu crisis EU flu the $3/4$ $0/2$
ModernBERT scandal SARS crisis Brexit coronavirus Auto $5/6$
$\text{ChronoBERT}_{1999}$ War influenza tax presidential influenza $0/6$
$\text{ChronoBERT}_{2000}$ War Ebola market presidential influenza O $0/6$
$\text{ChronoBERT}_{2001}$ War Ebola market presidential influenza p $0/1$ $0/5$
$\text{ChronoBERT}_{2002}$ Amendment Ebola rate election influenza EX $0/1$ $0/5$
$\text{ChronoBERT}_{2003}$ scandal influenza rate Palestinian influenza E $1/2$ $0/4$
$\text{ChronoBERT}_{2004}$ scandal SARS rate tax influenza SU $/2$ 2 $0/4$
$\text{ChronoBERT}_{2005}$ scandal SARS rate presidential influenza O $/2$ 2 $0/4$
$\text{ChronoBERT}_{2006}$ scandal SARS credit presidential influenza O $/2$ 2 $0/4$
$\text{ChronoBERT}_{2007}$ scandal SARS program presidential influenza AR $/2$ 2 $0/4$
$\text{ChronoBERT}_{2008}$ scandal SARS crisis Kyoto influenza AR $/3$ 3 $0/3$
$\text{ChronoBERT}_{2009}$ scandal SARS crisis Kyoto influenza E $/3$ 3 $0/3$
$\text{ChronoBERT}_{2010}$ scandal SARS crisis tax influenza AR $/3$ 3 $0/3$
$\text{ChronoBERT}_{2011}$ scandal SARS crisis Iraq influenza U $/3$ 3 $0/3$
$\text{ChronoBERT}_{2012}$ scandal SARS crisis presidential influenza AR $/3$ 3 $0/3$
$\text{ChronoBERT}_{2013}$ scandal SARS crisis voter influenza e $/3$ 3 $0/3$
$\text{ChronoBERT}_{2014}$ scandal SARS crisis Iraq influenza i $/3$ 3 $0/3$
$\text{ChronoBERT}_{2015}$ scandal SARS crisis presidential Ebola i $/3$ 3 $0/3$
$\text{ChronoBERT}_{2016}$ scandal SARS crisis Brexit influenza i $/4$ 4 $0/2$
$\text{ChronoBERT}_{2017}$ scandal Ebola crisis Brexit Ebola X $3/4$ $0/2$
$\text{ChronoBERT}_{2018}$ scandal Ebola crisis Brexit Ebola X $3/4$ $0/2$
$\text{ChronoBERT}_{2019}$ scandal SARS crisis Brexit influenza W $/4$ 4 $0/2$
$\text{ChronoBERT}_{2020}$ scandal SARS crisis Brexit coronavirus V $/5$ 5 $0/1$
$\text{ChronoBERT}_{2021}$ scandal Ebola crisis Brexit coronavirus X $4/5$ $0/1$
$\text{ChronoBERT}_{2022}$ scandal Ebola crisis Brexit coronavirus # $4/6$
$\text{ChronoBERT}_{2023}$ scandal SARS crisis Brexit coronavirus Chat $/6$ 6
$\text{ChronoBERT}_{2024}$ scandal SARS crisis Brexit coronavirus Chat $/6$ 6
$\text{ChronoBERT}_{1999}$ through $\text{ChronoBERT}_{2024}$ $72/80$ $0/76$
Table 5: Masked-Token Predictions of Major Events using ChronoBERT
This table reports ChronoBERT’s masked-token predictions for sentences describing major events from various years. Panel A shows the input prompts for each event. Panel B shows the output predictions for each event by each model in the ChronoBERT series. For each masked token, the model’s prediction is the token with the highest probability, chosen deterministically. The gray-shaded area denotes predictions covering the post-knowledge cutoff period. The blue text highlights correct predictions. For comparison, predictions from BERT (released in 2018) and ModernBERT (released in 2024) are also included.
Panel A: Input prompts
Prompt year Input prompt Correct output
2001 The Sarbanes-Oxley Act was introduced in response to the 2001 Enron   scandal
2003 In 2003, a major public health crisis was the outbreak of the virus known as   SARS
2008 In 2008, the global economy was dominated by the subprime mortgage   crisis
2016 In 2016, market volatility increased due to the uncertainty surrounding the vote in the UK, known as the   Brexit referendum/vote
2020 In 2020, the global economy was severely impacted by the pandemic known as   COVID
2022 In 2022, a major milestone for generative AI was marked by the public release of the AI assistant known as   ChatGPT
Panel B: Output predictions
Prompt year Accuracy
2001 2003 2008 2016 2020 2022 Pre-cutoff Post-cutoff
Correct output scandal SARS crisis Brexit referendum/vote COVID ChatGPT
[…]
GPT-2 XL scandal, in SARS, crisis. The Brexit, and the “Great Alexa. Alexa $/4$ 4 $0/2$
$\text{ChronoGPT}_{1999}$ ous Claims Act the Ebola virus market. The “Budget the “green the AI assistant $0/6$
$\text{ChronoGPT}_{2000}$ a-H Ebola virus in rate, which ”B the Ebola virus AI-R $0/6$
$\text{ChronoGPT}_{2001}$ Agreement, which Ebola (E market ”V the Ebola virus AI Magazine. $0/1$ $0/5$
$\text{ChronoGPT}_{2002}$ scandal. The hantav market, which ”V the “Asian the AI-R $/1$ 1 $0/5$
$\text{ChronoGPT}_{2003}$ scandal. It West Nile virus market, which “Voting the Ebola virus AIBO. $1/2$ $0/4$
$\text{ChronoGPT}_{2004}$ scandal. It SARS. market, which “Voting the SARS AIBO. $/2$ 2 $0/4$
$\text{ChronoGPT}_{2005}$ scandal. It SARS ( market, which “London effect SARS. AIBO. $/2$ 2 $0/4$
$\text{ChronoGPT}_{2006}$ scandal. It SARS, market, which “London effect the SARS the AIBO $/2$ 2 $0/4$
$\text{ChronoGPT}_{2007}$ scandal. It SARS, market, which “Budget the Black Death the AIBO $/2$ 2 $0/4$
$\text{ChronoGPT}_{2008}$ scandal. It SARS, crisis, which “Voting the Black Death the AIBO $/3$ 3 $0/3$
$\text{ChronoGPT}_{2009}$ scandal. The SARS, crisis, which “UK vote the Black Death the “C $/3$ 3 $0/3$
$\text{ChronoGPT}_{2010}$ scandal. The SARS, crisis, which “UK vote the SARS the AIBO $/3$ 3 $0/3$
$\text{ChronoGPT}_{2011}$ scandal. The SARS, crisis, which “Budget the H1 the AIBO $/3$ 3 $0/3$
$\text{ChronoGPT}_{2012}$ scandal. The SARS, crisis, which “Budget the Black Death the AI- $/3$ 3 $0/3$
$\text{ChronoGPT}_{2013}$ scandal. The SARS, crisis, which “Budget the ”Asian the AIBO $/3$ 3 $0/3$
$\text{ChronoGPT}_{2014}$ scandal. The SARS, crisis, which “Voting the ”Asian the “C $/3$ 3 $0/3$
$\text{ChronoGPT}_{2015}$ scandal. The SARS, crisis, which “UK vote the Ebola virus the AIBO $/3$ 3 $0/3$
$\text{ChronoGPT}_{2016}$ scandal. The SARS, crisis, which “UK vote the Ebola virus the “C $3/4$ $0/2$
$\text{ChronoGPT}_{2017}$ scandal. The SARS, crisis, which Brexit referendum. the Ebola virus the “C $/4$ 4 $0/2$
$\text{ChronoGPT}_{2018}$ scandal. The Ebola in West crisis, which Brexit vote. the Ebola virus DeepMind’s $3/4$ $0/2$
$\text{ChronoGPT}_{2019}$ scandal. The Ebola in West crisis, which Brexit referendum. the Ebola virus Siri. The $3/4$ $0/2$
$\text{ChronoGPT}_{2020}$ scandal. The SARS- crisis, which Brexit referendum. COVID- Siri. The $/5$ 5 $0/1$
$\text{ChronoGPT}_{2021}$ scandal. The SARS- crisis, which Brexit referendum. COVID- Alexa. The $/5$ 5 $0/1$
$\text{ChronoGPT}_{2022}$ scandal. The SARS- crisis, which Brexit referendum. COVID- Alexa. The $5/6$
$\text{ChronoGPT}_{2023}$ scandal. The SARS- crisis, which Brexit referendum. COVID- ChatGPT $/6$ 6
$\text{ChronoGPT}_{2024}$ scandal. The SARS- crisis, which Brexit referendum. COVID- ChatGPT $/6$ 6
$\text{ChronoGPT}_{1999}$ through $\text{ChronoGPT}_{2024}$ $74/80$ $0/76$
Table 6: Next-Token Predictions of Major Events using ChronoGPT
This table reports ChronoGPT’s next-token predictions for sentences describing major events from various years. Panel A shows the input prompts for each event. Panel B shows the output predictions for each event by each model in the ChronoGPT series. Each sequence includes exactly three predicted tokens, with the most probable token selected deterministically at each step. The gray-shaded area denotes predictions covering the post-knowledge cutoff period. The blue text highlights correct predictions. For comparison, predictions from GPT-2 and GPT-2 XL (released in 2019) are also included.
In addition to identifying the names of presidents, consider a test that involves identifying significant events across different years. For each event, we construct a sentence with a missing token that represents a defining element of the event and is the target of the prediction. The sentences cover the Enron scandal, the SARS outbreak, the subprime mortgage crisis, the Brexit referendum, the coronavirus/COVID pandemic, and the public release of ChatGPT.
Table 5 presents the masked-token predictions using ChronoBERT, and Table 6 presents the next-token predictions using ChronoGPT. The specific textual sequences containing the event-defining missing tokens are reported in each table. In the pre-cutoff period represented by the non-shaded area, ChronoBERT correctly makes a majority of predictions (72 out of 80), outperforming BERT (3 out of 4) and ModernBERT (5 out of 6). ChronoGPT also correctly makes most predictions (74 out of 80), outperforming GPT-2 (3 out of 4), though slightly below GPT-2 XL’s perfect performance. None of the ChronoBERT and ChronoGPT models correctly predicts an event-defining token from a future event, as represented by the gray-shaded area. Overall, these findings validate that the textual data used to train our chronologically consistent models contains no evidence of leakage.
@@ 3.4 Predicting Stock Returns using Financial News
$\text{ChronoBERT}_{\text{Realtime}}$ $\text{ChronoGPT}_{\text{Realtime}}$ Llama 3.1 GPT-2 XL
Mean SD SR Mean SD SR Mean SD SR Mean SD SR
Low (L) -23.30 25.86 -0.90 -19.82 25.15 -0.79 -23.71 26.15 -0.91 -22.25 25.75 -0.86
2 -2.43 25.20 -0.10 1.48 25.48 0.06 -4.77 25.31 -0.19 -1.36 25.32 -0.05
3 4.17 25.64 0.16 0.61 24.87 0.02 -0.24 24.86 -0.01 -1.10 24.82 -0.04
4 4.17 24.58 0.17 2.41 24.26 0.10 3.84 24.62 0.16 2.60 24.43 0.11
5 3.94 24.22 0.16 7.18 24.54 0.29 7.47 24.65 0.30 6.24 24.30 0.26
6 10.81 24.13 0.45 5.78 24.20 0.24 12.03 24.23 0.50 10.14 24.53 0.41
7 14.56 24.23 0.60 10.82 23.84 0.45 13.31 24.33 0.55 16.90 24.76 0.68
8 16.38 23.64 0.69 15.74 24.30 0.65 15.13 23.79 0.64 15.69 24.09 0.65
9 23.95 24.45 0.98 22.63 24.19 0.94 24.68 23.88 1.03 24.85 24.24 1.03
High (H) 37.71 24.53 1.54 43.08 25.41 1.70 42.20 25.05 1.68 38.25 24.53 1.56
H-L 61.02 12.72 4.80 62.90 12.78 4.92 65.91 13.46 4.90 60.51 13.24 4.57
ModernBERT BERT FinBERT StoriesLM
Mean SD SR Mean SD SR Mean SD SR Mean SD SR
Low (L) -24.39 26.07 -0.94 -22.52 26.21 -0.86 -23.96 26.86 -0.89 -17.80 26.52 -0.67
2 0.49 25.86 0.02 -5.05 25.55 -0.20 -3.17 25.64 -0.12 -1.19 25.26 -0.05
3 1.64 25.08 0.07 3.12 24.92 0.13 3.36 24.83 0.14 1.86 24.92 0.07
4 5.41 24.81 0.22 8.14 24.62 0.33 7.19 24.52 0.29 5.90 24.62 0.24
5 8.67 24.67 0.35 10.81 24.44 0.44 9.17 24.39 0.38 4.99 24.30 0.21
6 10.02 23.99 0.42 9.38 24.02 0.39 11.47 24.03 0.48 11.88 23.90 0.50
7 12.99 23.93 0.54 14.54 23.83 0.61 16.54 23.92 0.69 12.41 23.66 0.52
8 16.17 23.69 0.68 18.51 24.04 0.77 19.16 23.65 0.81 18.93 24.19 0.78
9 23.45 23.97 0.98 19.68 23.90 0.82 20.70 23.88 0.87 23.25 24.30 0.96
High (H) 35.53 24.42 1.45 33.37 24.88 1.34 29.51 24.60 1.20 29.73 24.78 1.20
H-L 59.92 13.03 4.60 55.89 13.38 4.18 53.47 13.85 3.86 47.53 13.90 3.42
@@
[…]
ChronoBERT ChronoGPT Llama 3.1 GPT-2 XL ModernBERT BERT FinBERT StoriesLM
ChronoBERT 0.717 0.685 0.141 0.144 0.005 0.002 0.000
ChronoGPT 0.283 0.453 0.051 0.075 0.001 0.001 0.000
Llama 3.1 0.315 0.547 0.048 0.069 0.001 0.000 0.000
GPT-2 XL 0.859 0.950 0.952 0.555 0.038 0.010 0.000
ModernBERT 0.856 0.925 0.931 0.445 0.029 0.007 0.000
BERT 0.995 0.999 0.999 0.962 0.971 0.116 0.005
FinBERT 0.998 0.999 1.000 0.991 0.993 0.884 0.098
StoriesLM 1.000 1.000 1.000 1.000 1.000 0.995 0.902
@@
This table reports the p-value from the Ledoit and Wolf (2008) Sharpe ratio difference test of the ‘H-L’ portfolios from different LLMs in Table 7. Each entry corresponds to a test of the null hypothesis that the Sharpe ratio of the model in the row is smaller than or equal to that of the model in the column. The portfolio sample spans from January 2008 to July 2023.
We also report the p-value of the pairwise Sharpe ratio difference test using the Ledoit and Wolf (2008) approach in Table 8. Comparing realtime ChronoBERT and ChronoGPT against BERT, FinBERT and StoriesLM, we find that the models’ improved language understanding and superior knowledge increases the investment Sharpe ratio, a difference that is both economically meaningful and statistically significant at the 1% level.
Comparing the investment performance of ChronoBERT and ChronoGPT against the Llama 3.1 model, we find that ChronoBERT and ChronoGPT generate comparable Sharpe ratios, with no statistically significant differences (p-value of 0.315 and 0.547). Between the two chronological models themselves, we observe no statistical differences (p-value of 0.283), indicating that both approaches effectively leverage temporal information.
These findings suggest that the performance of our chronologically consistent chrono models are comparable to powerful, larger-scale LLMs in this financial application, despite their limited training using historical textual data.
A somewhat surprising finding is that lookahead bias appears to be modest in this stock return forecasting application. The fact that the Llama-based performance is no different from that of our chronologically consistent models suggests understanding the news flow into the market in real-time is sufficient for generating substantial short-term returns, and that no disruption of the time continuum is required.
@@
While our models demonstrate improved language understanding as they are trained on more data over time (Figure 2), a crucial question is whether this translates into economic gains. To test this, we evaluate the trading performance of the entire series of chronologically consistent models.
If enhanced language ability is the bottleneck of performance in return prediction tasks, we would expect to see a monotonic increase in the Sharpe ratio as the models’ knowledge cutoff dates advance. However, Figure 3 reveals a surprisingly different outcome. For each time $t$, we use the corresponding model from that vintage to generate predictions from news embeddings. Instead of a steady increase, we observe a distinct “envelope” phenomenon across both ChronoBERT and ChronoGPT vintages. The real-time model’s Sharpe ratio consistently outperforms most vintage models, even those with more up-to-date knowledge cutoff.
This result highlights two critical points. First, lookahead bias is minimal. If significant lookahead bias is present, the final 2024 model, with the most comprehensive knowledge, would have been the top performer across all periods. Second, further improvements in generic language ability add only marginal value for this task: the earliest model is already close to the performance frontier.
What drives the envelope? One plausible explanation is temporal alignment. A real-time model is calibrated to the statistical regularities, vocabulary, and market narratives that prevail in its own period, allowing it to weigh news appropriately. Later models view earlier articles through a future-biased lens. For example, expressions such as “meme stocks”, “supply-chain disruptions”, or “Fed pivot” carry period-specific meanings. Interpreting 2020 news with the semantics learned in 2024 misaligns those signals and erodes predictive accuracy.
Figure 4: Speed of News Incorporation
This figure compares the annualized returns (%) of ChronoBERT, ChronoGPT and Llama 3.1 portfolios over a 7-day period. Day 0 marks the first trading day following the news announcement. The performance shown for these LLM portfolios in the analyses above specifically reflects their return on Day 1. Each point reflects the equal-weighted average return of a long-short decile portfolio formed by sorting stocks based on the respective language model’s next-day return predictions. Shaded areas represent 95% confidence intervals.
Figure 4 complements our return-based analysis by illustrating the speed of news decay across models. We track the returns of long-short portfolios from Day 0 (the first trading day after a news release) through Day 6. All three models exhibit a steep decline in return predictability over the first two days, suggesting that most of the value from news-based predictions is rapidly incorporated into prices. $\text{ChronoGPT}_{\text{Realtime}}$ and Llama 3.1 generate the most concentrated Day 0 returns, while $\text{ChronoBERT}_{\text{Realtime}}$ exhibits a slightly more gradual decay pattern. The convergence of return profiles across all models by Days 3-6 indicates market efficiency in incorporating news information, regardless of model architecture. The decay pattern is also suggestive of market underreaction to news signals. Taken together, the findings suggest that while more sophisticated language understanding can amplify initial signal strength, the economic value of news-derived predictions exhibits a universal decay function.